solve for x: 3^5x+1=5
Answer by lwsshak3(6491)

solve for x: 3^5x+1=5
3^5x+1=5
3^5x=5-1=4
take log of both sides
5xlog3=log4
5x=log4/log3
x=(log4/log3)/5
using calculator
x≈.2524